# Semi-Supervised Self-Learning Enhanced Music Emotion Recognition

### v2

@@
At the beginning of each epoch, the training set is partitioned into a clean set and a labeled noisy set using a two-component Gaussian mixture model (GMM) by the cross-entropy loss value for each training sample. Then the expectation-maximization (EM) algorithms are leveraged to fit the GMM to the observation. Then, the semi-supervised learning method is used to treat the clean set as the labeled set $\mathcal{X}$ while the noisy set is the unlabeled set $\mathcal{U}$. We erase the label of the $\mathcal{U}$ set, use the predicted value of the model as the soft label, and then use a sharpening algorithm to get the pseudo soft label. Then we obtain the total loss using mixup and consistency regularization.
@@
In a k-class classification problem, the training data with n training samples denoted as $X=[x_{1},\ldots,x_{n}]$, along with their corresponding ground-truth labels $Y=[y_{1},\ldots,y_{n}]$. Each $y_{i}$ represents the true class label and is represented as a k-dimensional one-hot vector. Classification problems on clean label datasets are often defined as:
@@
where ${\theta}$ represents the model parameters, and $\mathcal{L}$ represents the loss objective function. The cross-entropy loss function in a classification problem is defined as follows:
@@
where $f$ denotes the output of the neural network classifier, i.e. the softmax layer. But when we use the above formulas (1) and (2) to train on the noisy label dataset, severe overfitting will occur.
@@
DNNs have been observed to prioritize learning from simple and logically consistent samples in the presence of noisy labels, resulting in reduced loss for these samples, particularly in the early stages of training [12]. This phenomenon suggests that the loss distributions of clean and noisy samples during training can be approximated by two Gaussian distributions, with the clean samples having a smaller mean loss. Leveraging this training characteristic, we employ a GMM to differentiate between noisy and clean samples by utilizing the per-sample loss as input. The probability density function of the loss K-component mixture model is defined as Eq. (3):
@@
where $\lambda_{k}$ indicates the mixing coefficients for the combination of each probability density function $p(l|k)$. Regarding this case, we can utilize a two-component GMM to model the distribution of clean and noisy samples. We utilize a two-component GMM and then feed the cross-entropy loss $\mathcal{L}$ of every training data to the GMM. Then, to estimate the parameters of the GMM we apply the EM algorithm. We define the posterior probability $\omega_{i}$ as the probability that the $i_{th}$ data sample belongs to the Gaussian component with a smaller mean, given its loss $l_{i}$. By setting a global threshold $\tau$ for the probability $\omega$ of all training samples’ labels to be clean, we can split the expanded segment-level dataset $\mathcal{D}$ into two parts: set $\mathcal{X}$ and set $\mathcal{U}$. The labels in set $\mathcal{X}$ have more possibilities to be correct, and thus, they will be used as the labeled set. On the other hand, the labels in set $\mathcal{U}$ are more likely to be incorrect, and will be erased in the subsequent steps. Therefore, set $\mathcal{U}$ will be used as the unlabeled set.
@@
We first preprocess the set $\mathcal{U}$. For the set $\mathcal{U}$, the initial labels are likely to be incorrect and have been erased. As a result, we generate pseudo soft labels $\hat{y}$ by sharpening the predicted distribution of the model by Eq. (4).
@@
where $S(\cdot)$ indicates the temperature sharpening function often used in pseudo labeling, and $T$ represents the temperature coefficient. And the $\mathcal{\hat{U}}$ will be generated then as Eq. (5).
@@
As [13] has shown the mixup technology can eliminate confirmation bias to a certain extent. This technique involves training on convex combinations of pairs of samples, denoted as $x_{p}$ and $x_{q}$, along with their corresponding labels $y_{p}$ and $y_{q}$ shown as follows:
@@
where $\delta$ is drawn from a beta distribution randomly. This integration introduces regularization to encourage the network to exhibit linear behavior between samples, thereby reducing fluctuations in distant regions. In terms of label noise, mixup offers a strategy to merge clean and noisy samples, resulting in a more representative loss that guides the training procedure.
@@
Eq. (6) can be regarded in the loss as $l=\delta l_{p}+(1-\delta)l_{q}$, and the standard cross-entropy loss is leveraged for semi-supervised learning part by Eq. (7):
@@
Therefore, the total loss is as Eq. (9):
@@
A reliable segment-level classifier serves as a prerequisite for clip-level classification. This allows us to utilize machine learning algorithms to handle structured features obtained from statistical properties of the segment probability distributions. The probability of the $k_{th}$ emotion for segment $s$ is defined as $P_{s}(E_{k})$ in Eq. (10):
@@
where $C$ represents the set of all segments in a single song. Features $f_{1}^{k}$, $f_{2}^{k}$ and $f_{3}^{k}$ correspond to the maximum value of the segment-level probability of the $k_{th}$ emotion in the song, minimum, and average values, respectively. The features $f_{3}^{k}\sim f_{5}^{k}$ correspond to the $k_{th}$ emotion’s three quartiles in the song respectively. The feature $f_{7}^{k}$ represents the percentage of segments with a high probability of sentiment k. The feature exhibits low sensitivity to a threshold, allowing for empirical selection. This aggregation step produces a feature representation of dimension $7\times K$ for each song. With this set of song-level feature representations, we can train a secondary classifier that is relatively simple in nature to make decisions at the song level.

### v3

@@
At the beginning of each epoch, the training set is partitioned into a clean set and a labeled noisy set using a two-component Gaussian mixture model (GMM) by the cross-entropy (CE) loss value for each training sample. Then the expectation-maximization (EM) algorithms are leveraged to fit the GMM to the observation. Then, the semi-supervised learning method is used to treat the clean set as the labeled set $\mathcal{X}$ while the noisy set is the unlabeled set $\mathcal{U}$. We erase the label of the $\mathcal{U}$ set, use the predicted value of the model as the soft label, and then use a sharpening algorithm to get the pseudo soft label. Then we obtain the total loss using mixup and consistency regularization.
@@
In a k-class classification problem, the training data with n training samples denoted as $X=[x_{1},\ldots,x_{n}]$, along with their corresponding ground-truth labels $Y=[y_{1},\ldots,y_{n}]$. Each $y_{i}$ represents the true class label and is represented as a k-dimensional one-hot vector. Classification problems on clean label datasets are often defined as Eq. (1), where ${\theta}$ represents the model parameters, and $\mathcal{L}$ represents the loss objective function.
@@
The CE loss function in a classification problem is defined as Eq. (2), where $f$ denotes the output of the neural network classifier, i.e. the softmax layer.
@@
But when we use the above formulas (1) and (2) to train on the noisy label dataset, severe overfitting will occur.
@@
DNNs have been observed to prioritize learning from simple and logically consistent samples in the presence of noisy labels, resulting in reduced loss for these samples, particularly in the early stages of training [12]. This phenomenon suggests that the loss distributions of clean and noisy samples during training can be approximated by two Gaussian distributions, with the clean samples having a smaller mean loss. Leveraging this training characteristic, we employ a GMM to differentiate between noisy and clean samples by utilizing the per-sample loss as input. The probability density function of the loss K-component mixture model is defined as Eq. (3), where $\lambda_{k}$ indicates the mixing coefficients for the combination of each probability density function $p(l|k)$.
@@
Regarding this case, we can utilize a two-component GMM to model the distribution of clean and noisy samples. We utilize a two-component GMM and then feed the CE loss $\mathcal{L}$ of every training data to the GMM. Then, to estimate the parameters of the GMM we apply the EM algorithm. We define the posterior probability $\omega_{i}$ as the probability that the $i_{th}$ data sample belongs to the Gaussian component with a smaller mean, given its loss $l_{i}$. By setting a global threshold $\tau$ for the probability $\omega$ of all training samples’ labels to be clean, we can split the expanded segment-level dataset $\mathcal{D}$ into two parts: set $\mathcal{X}$ and set $\mathcal{U}$. The labels in set $\mathcal{X}$ have more possibilities to be correct, and thus, they will be used as the labeled set. On the other hand, the labels in set $\mathcal{U}$ are more likely to be incorrect, and will be erased in the subsequent steps. Therefore, set $\mathcal{U}$ will be used as the unlabeled set.
@@
We first preprocess the set $\mathcal{U}$. For the set $\mathcal{U}$, the initial labels are likely to be incorrect and have been erased. As a result, we generate pseudo soft labels $\hat{y}$ by sharpening the predicted distribution of the model by Eq. (4), where $S(\cdot)$ indicates the temperature sharpening function often used in pseudo labeling, and $T$ represents the temperature coefficient.
@@
And the $\mathcal{\hat{U}}$ will be generated then as Eq. (5).
@@
As [13] has shown the mixup technology can eliminate confirmation bias to a certain extent. This technique involves training on convex combinations of pairs of samples, denoted as $x_{p}$ and $x_{q}$, along with their corresponding labels $y_{p}$ and $y_{q}$ shown as Eq. (6), where $\delta$ is drawn from a beta distribution randomly.
@@
This integration introduces regularization to encourage the network to exhibit linear behavior between samples, thereby reducing fluctuations in distant regions. In terms of label noise, mixup offers a strategy to merge clean and noisy samples, resulting in a more representative loss that guides the training procedure.
@@
Eq. (6) can be regarded in the loss as $l=\delta l_{p}+(1-\delta)l_{q}$, and the standard CE loss is leveraged for semi-supervised learning part by Eq. (7):
@@
Therefore, the total loss is as Eq. (9), where $\lambda$ is the hyperparameter to control the weight.
@@
A reliable segment-level classifier serves as a prerequisite for clip-level classification. This allows us to utilize machine learning algorithms to handle structured features obtained from statistical properties of the segment probability distributions. The probability of the $k_{th}$ emotion for segment $s$ is defined as $P_{s}(E_{k})$ in Eq. (10), where $C$ represents the set of all segments in a single song.
@@
Features $f_{1}^{k}$, $f_{2}^{k}$ and $f_{3}^{k}$ correspond to the maximum value of the segment-level probability of the $k_{th}$ emotion in the song, minimum, and average values, respectively. The features $f_{3}^{k}\sim f_{5}^{k}$ correspond to the $k_{th}$ emotion’s three quartiles in the song respectively. The feature $f_{7}^{k}$ represents the percentage of segments with a high probability of sentiment k. The feature exhibits low sensitivity to a threshold, allowing for empirical selection. This aggregation step produces a feature representation of dimension $7\times K$ for each song. With this set of song-level feature representations, we can train a secondary classifier that is relatively simple in nature to make decisions at the song level.